OpenStudy (anonymous):

The solute potassium chloride (KCl) is dissolved in the solvent water (H2O) to produce a homogeneous solution. The solution contains 15.0% of KCl by mass. What mass of water is contained in 200. grams of this solution?

OpenStudy (aaronq):

if you have a 15% by mass solution of KCl, it means you have 15 g of KCl in 100 g of solution (meaning you have 85 grams of water in 100 grams of solution) 100g-15g=85 g so in 200 g of solution you have 30 grams of KCl, right?

OpenStudy (jebonna):

Find out 15% of 200g 200/100x15 = 30g of KCl 200g-30g = 170g of H2O So surely it would be 170g of water by mass and 30g of KCl

OpenStudy (anonymous):

Oh. That was a dumb question. Thanks, guys.